Kingston anticipates total sales for June and July of $380,000 and $338,000, respectively. Cash sales are normally 65% of total sales. Of the credit sales, 20% are collected in the same month as the sale, 65% are collected during the first month after the sale, and the remaining 15% are collected in the second month after the sale. Determine the amount of accounts receivable reported on the companys budgeted balance sheet as of July 31. As of July 31 Sales month Total Sales Credit SalesUncollected Uncollected Amount June July Total $ 380,000 338,000
Answer #1

Requirement--

As of Jul 31
Sales month Total Sales Credit sales % still uncollected Amount uncollected
June $380,000 $133,000 15% $19,950
Jul $338,000 $118,300 80% $94,640
Total $114,590

--Working

As of Jul 31
Sales month Total Sales Credit sales % still uncollected Amount uncollected
June 380000 =380000*35% 0.15 =133000*15%
Jul 338000 =338000*35% 0.8 =118300*80%
Total =19950+94640